# Unified Flavor:

### 11 Mar 2026 v1

@@ I Introduction
It is worth emphasizing that the numerator of a ninths exponent is not the chain length: it is the sum of discrete charge differences along the chain. If one simply assigned the flavon a charge of $1/9$, then integrating out $n$ messengers would give $\epsilon^{n/9}$ with $n$ equal to the number of links. That would require 18 messengers to produce $|V_{cb}|\sim\epsilon^{2}$, or 27 for $|V_{ub}|\sim\epsilon^{3}$—hardly economical. In the $B$-lattice framework the flavon carries charge 1 under $Z_{9}$, but the messenger sites carry different $Z_{9}$ charges, so the charge gap $h_{k}$ at each nearest-neighbor hop need not be unity. For instance, a three-site chain with hop charges $(h_{1},h_{2},h_{3})=(1,2,4)\;\mathrm{mod}\;9$ yields a total exponent $(1+2+4)/9=7/9$ from only three links. Because the set $\{1,2,4\}$ generates all residues mod 9, different Yukawa entries—corresponding to different endpoint dressings on the same short chain—access the full ninths lattice $\{n/9\}$ without requiring long chains. The denominator 9 thus reflects the order of the discrete gauge group, while the numerator encodes the charge assignment pattern at the chain sites.
@@ II.1 Symmetry structure
where $Z_{N}$ (with $N=9$ or $18$) enforces the ninths-quantized suppression of FN operators, and $Z_{2}^{\rm(NN)}$ forbids nonlocal (non-nearest-neighbor) chain couplings. The discrete gauge anomaly cancellation conditions of Ibáñez and Ross [4] are satisfied generation by generation for $Z_{9}$ or $Z_{18}$.
@@ III.1 Down-type vectorlike chain
giving a discrete hop set $\{1,2,4\}/9$. Note that the number of flavon insertions at each hop equals the $Z_{9}$ charge difference, ensuring gauge invariance. The total chain suppression is $\epsilon^{(1+2+4)/9}=\epsilon^{7/9}$.
@@
Figure 1: Schematic of the four-site down-type VLQ chain. Solid arrows denote nearest-neighbor hops with the indicated $\epsilon^{h_{a}/9}$ suppression; dashed arrows show the endpoint couplings to SM fields. The $Z_{9}$ charges are shown below each site. The effective Yukawa exponent for entry $(i,j)$ is $(\Sigma+A_{i}+B_{j})/9$, an exact result of the chain-inversion theorem (Sec. IV), not a perturbative approximation.
@@ III.3 Up-type chain
Figure 2: Combined down-type (blue, upper) and up-type (red, lower) VLQ chains. Both chains share the same internal hop structure $\{1,2,4\}/9$ and the same left-handed endpoint dressings $A_{i}=(27,18,0)$, enforced by SU(2)L gauge invariance acting on the shared quark doublet $q_{L,i}$. The right-handed dressings differ: $B_{j}^{d}=(10,3,0)$ for the down sector and $B_{j}^{u}=(37,12,0)$ for the up sector. The $\epsilon^{h/9}$ suppressions and the resulting Yukawa exponents are exact consequences of the upper-triangular chain mass matrix (Sec. IV.5); they do not rely on $\kappa_{a}\epsilon^{h_{a}/9}\ll 1$. The CKM matrix arises from the mismatch between the left-handed rotations $U_{dL}$ and $U_{uL}$ that diagonalize the two Yukawa textures.
@@ VI.7 Jarlskog invariant scaling
$J\sim\epsilon^{8/9}\cdot\epsilon^{17/9}\cdot\epsilon^{10/3}\cdot\sin\delta\sim\epsilon^{19/3}\,\sin\delta\simeq 2.5\times 10^{-5}\sin\delta,$ (51)
@@ VII.2 Genericity of $\mathcal{O}(1)$ CP violation
Phase counting. The standard phase-counting theorem for three generations gives (n1)(n2)/2=1(n-1)(n-2)/2=1 physical CP-violating phase [24]. The multi-messenger phases $\phi_{f}$, $\psi_{f}$ are not additional parameters: they are specific combinations of the UV coupling phases that survive after all field rephasing. The chain framework therefore has the same number of irreducible phase inputs as the SM—one. Its advantage is that the lattice exponent algebra determines the scaling $J\sim\epsilon^{19/3}\sin\delta$ structurally, while the dense ninths spacing ensures that the remaining free parameter $\sin\delta$ is generically $\mathcal{O}(1)$.
@@ XII.1 The charged-lepton GMR and $\epsilon^{3/2}$
and numerically $\epsilon^{3/2}\simeq 0.0806$—an exact match. The exponent $3/2=Q(L_{1})+Q(L_{2})=1+\tfrac{1}{2}$ is the sum of the first- and second-generation lepton-doublet charges, so $\beta$ is determined by the $Z_{18}$ discrete gauge charges.
@@ XII.4 Dual mass representations of CKM elements
These identities are not numerological coincidences: they are algebraic consequences of the fact that all six quark masses are determined by a single parameter $B$ and the discrete $Z_{9}$ charges. In the $B$-lattice, the masses scale as $(u,c,t)=t\,(\epsilon^{64/9},\,\epsilon^{10/3},\,1)$ and $(d,s,b)=b\,(\epsilon^{37/9},\,\epsilon^{7/3},\,1)$, so any ratio of mass monomials reduces to a rational power of $\epsilon$. A complete catalog of the resulting dual CKM representations and geometric mean relations will be presented in a forthcoming companion paper.
@@ XIII Discrete Anomaly Cancellation
XIII.1 $Z_{18}$ charge assignments
The complete $Z_{18}$ charge table follows from the $B$-lattice FN charges determined in Secs. V and XI [5, 21], expressed as integer charges $q_{18}\equiv 18\,Q$:
@@ XIII.1 $Z_{18}$ charge assignments
Here $N_{i}$ are right-handed (sterile) neutrinos and $H$ is the SM Higgs doublet. The third generation of all SM fields carries zero charge, so the anomaly sums receive contributions only from the first two generations. The $Z_{9}$ subgroup (denominator-$9$ lattice) controls quark-sector exponents, while the full $Z_{18}$ is required to accommodate the half-integer lepton-doublet charge $Q(L_{2})=\tfrac{1}{2}$.
@@ XIII.2 Neutrino masses from the charge table
The $Z_{18}$ charges determine the neutrino mass texture through two complementary mechanisms.
@@
The dimension-five operator $L_{i}L_{j}HH/\Lambda$ carries $Z_{18}$ charge $Q(L_{i})+Q(L_{j})$, so the required number of flavon insertions gives the exponent matrix $p^{\nu}_{ij}=Q(L_{i})+Q(L_{j})$ of Eq. (79).
@@
The vanishing $Z_{18}$ charge of $N_{i}$ has two further consequences: the right-handed neutrinos are SM singlets with zero discrete charge, so they do not contribute to any of the anomaly sums (106)–(108); and their Majorana mass $M_{N}\sim\Lambda$ is unsuppressed by the flavon, placing the seesaw scale at $\Lambda\sim 5\times 10^{11}$ GeV, consistent with the axion decay constant $f_{a}$ discussed in Sec. XIV.
@@ XIII.3 Green–Schwarz cancellation and the axion
The SM fermion charges above do not by themselves satisfy Eqs. (106)–(108); the residual anomalies $A_{a}\not\equiv 0\pmod{18}$ are nonzero. This is a generic feature of FN models and is resolved by the Green–Schwarz (GS) mechanism: the $Z_{18}$ discrete gauge symmetry descends from a continuous $U(1)_{\rm FN}$ that is spontaneously broken by the flavon VEV $\langle\Phi\rangle$. The $U(1)_{\rm FN}$ mixed anomalies are cancelled by a GS axion, and this cancellation persists in the discrete $Z_{N}$ remnant [4, 38].
The connection to the strong CP problem is immediate: the same PQ axion that solves the strong CP problem (Sec. XIV) can serve as the GS field that cancels the discrete gauge anomalies. This identification is natural in the “Flavor in Ninths” framework [21], where the $Z_{18}$ discrete gauge symmetry simultaneously
@@
The VLQ chain fields ($D_{a}+\bar{D}_{a}$, $U_{a}+\bar{U}_{a}$) are vectorlike under all gauge groups including $Z_{18}$ and therefore do not contribute to the anomaly sums.
@@ XIII.4 Spectator fermion alternative
If the GS mechanism is not invoked, anomaly cancellation can alternatively be achieved by introducing spectator fermions at a high scale $\Lambda\gg M_{\rm VLQ}$ that carry appropriate $Z_{18}$ charges but are SM singlets. Such spectators are common in UV completions of FN models and decouple from low-energy physics. In either case, the anomaly-free $Z_{18}$ is a legitimate discrete gauge symmetry, and its protection of the axion quality and the lattice exponent structure is on firm theoretical footing.
@@ XIV Connection to the Axion Quality Problem
The “Flavor in Ninths” paper [21] demonstrated that the same discrete $Z_{N}$ symmetry enforcing the ninths lattice can simultaneously address the Peccei–Quinn (PQ) axion quality problem [39]. The axion quality problem—the requirement that Planck-suppressed operators violating the global $U(1)_{\rm PQ}$ symmetry be forbidden to sufficiently high dimension—was identified by Kamionkowski and March-Russell [40], Holman et al. [41], and Barr and Seckel [42], and has been studied quantitatively in the context of the string landscape by Baer, Barger, and Sengupta [43] and in gauge-protected axion models by Bhattiprolu and Martin [44]. In the UF construction, the FN flavon $\Phi$ is identified with the PQ scalar, and the $Z_{18}$ discrete gauge symmetry forbids all PQ-violating Planck-suppressed operators below dimension $9$.
@@
This connection carries over to the chain framework: the $Z_{9}$ (or $Z_{18}$) symmetry that enforces nearest-neighbor locality along the VLQ chain is the same symmetry that protects the axion potential. The chain construction therefore inherits the axion-quality protection automatically, without requiring any additional symmetry.
@@
The axion in this mass window ($f_{a}\sim 10^{11}$ GeV) is a viable dark matter candidate. The continued absence of weakly interacting massive particle (WIMP) signals in direct detection experiments [47, 48, 49] suggests that WIMPs—if they were ever thermally produced—may have decayed away via small symmetry-violating couplings, leaving axions as the dominant component of dark matter today [50, 51]. In the UF framework this outcome is natural: the same $Z_{18}$ discrete gauge symmetry that protects the axion quality also forbids the lowest-dimension operators that would destabilize the lightest supersymmetric particle in SUSY extensions, providing a unified explanation for both the flavor hierarchy and the dark matter composition.
@@ XVI Discussion
    The discrete anomaly of $Z_{18}$ is cancelled by the Green–Schwarz mechanism, identifying the GS axion with the PQ axion that solves the strong CP problem. The specific group order $N=18$ is not a free choice: it is the minimal $Z_{N}$ that simultaneously accommodates the half-integer lepton doublet charge $Q(L_{2})=\tfrac{1}{2}$, the ninths-quantized quark exponents, and the axion quality requirement that PQ-violating operators be forbidden below dimension $N/2=9$.
@@
Axion quality $Z_{18}$ protected Haloscopes
@@ XVII Conclusions
Unified Flavor provides the dynamical ultraviolet completion of the $B$-lattice hierarchy program. A single flavon $\Phi$ with $\epsilon=1/B=14/75$ enforces discrete ninths-quantized Yukawa exponents, while TeV-scale nearest-neighbor VLQ chains generate the effective Yukawa couplings as algebraic path sums. Each effective Yukawa amplitude factorizes into entry, chain-propagation, and exit factors controlled by the discrete $Z_{9}$ charges of the SM fields, with the mass hierarchy encoded in the exponents $\epsilon^{Q(Q_{i})+Q(f^{c}_{j})}$. A multi-messenger structure—coherent contributions from several chain configurations—generates $\mathcal{O}(1)$ complex coefficients whose phases are the physical origin of CKM CP violation.
@@
Unified Flavor completes the $B$-lattice series by demonstrating that the phenomenologically successful lattice exponent algebra emerges from a definite symmetry structure and field content, rendering the fermion mass hierarchy a structural consequence of the UV theory rather than a parametric accident.
@@ D.3 Physical interpretation
(iii) One irreducible phase. The standard phase-counting theorem gives (n1)(n2)/2=1(n-1)(n-2)/2=1 physical CP phase for three generations. The multi-messenger phases $\phi_{f}$, $\psi_{f}$ are specific combinations of the UV Yukawa coupling phases that survive after all field rephasing; they do not represent additional parametric freedom. The framework’s structural claim is that the Jarlskog scaling $J\sim\epsilon^{19/3}\sin\delta$ is fixed by the lattice exponents, while the remaining free parameter $\sin\delta$ is generically $\mathcal{O}(1)$—it is not tuned to be either large or small.

### 13 Mar 2026 v2

@@ I Introduction
It is worth emphasizing that the numerator of a ninths exponent is not the chain length: it is the sum of discrete charge differences along the chain. If one simply assigned the flavon a charge of $1/9$, then integrating out $n$ messengers would give $\epsilon^{n/9}$ with $n$ equal to the number of links. That would require 17 messengers to produce $|V_{cb}|\sim\epsilon^{17/9}$, or 30 for $|V_{ub}|\sim\epsilon^{10/3}$—hardly economical. In the $B$-lattice framework the flavon carries charge 1 under $\mathbb{Z}_{9}$, but the messenger sites carry different $\mathbb{Z}_{9}$ charges, so the charge gap $h_{k}$ at each nearest-neighbor hop need not be unity. For instance, a three-site chain with hop charges $(h_{1},h_{2},h_{3})=(1,2,4)\;\mathrm{mod}\;9$ yields a total exponent $(1+2+4)/9=7/9$ from only three links. Because the set $\{1,2,4\}$ generates all residues mod 9, different Yukawa entries—corresponding to different endpoint dressings on the same short chain—access the full ninths lattice $\{n/9\}$ without requiring long chains. The denominator 9 thus reflects the order of the discrete gauge group, while the numerator encodes the charge assignment pattern at the chain sites.
@@ II.1 Symmetry structure
where $Z_{N}$ (with $N=9$ or $18$) enforces the ninths-quantized suppression of FN operators, and $Z_{2}^{\rm(NN)}$ forbids nonlocal (non-nearest-neighbor) chain couplings. The discrete gauge anomaly cancellation conditions of Ibáñez and Ross [4] are satisfied generation by generation for $\mathbb{Z}_{9}$ or $\mathbb{Z}_{18}$.
@@ III.1 Down-type vectorlike chain
giving a discrete hop set $\{1,2,4\}/9$. Note that the number of flavon insertions at each hop equals the $\mathbb{Z}_{9}$ charge difference, ensuring gauge invariance. The total chain suppression is $\epsilon^{(1+2+4)/9}=\epsilon^{7/9}$.
@@
Figure 1: Schematic of the four-site down-type VLQ chain. Solid arrows denote nearest-neighbor hops with the indicated $\epsilon^{h_{a}/9}$ suppression; dashed arrows show the endpoint couplings to SM fields. The $\mathbb{Z}_{9}$ charges are shown below each site. The effective Yukawa exponent for entry $(i,j)$ is $(\Sigma+A_{i}+B_{j})/9$, an exact result of the chain-inversion theorem (Sec. IV), not a perturbative approximation.
@@ III.3 Up-type chain
Figure 2: Combined down-type (blue, upper) and up-type (red, lower) VLQ chains. Both chains share the same internal hop structure $\{1,2,4\}/9$ and the same left-handed endpoint dressings $A_{i}=(27,18,0)$ [Eq. (32)], enforced by SU(2)L gauge invariance acting on the shared quark doublet $q_{L,i}$. The right-handed dressings differ [Eq. (33)]: $B_{j}^{d}=(10,3,0)$ for the down sector and $B_{j}^{u}=(37,12,0)$ for the up sector. The $\epsilon^{h/9}$ suppressions and the resulting Yukawa exponents are exact consequences of the upper-triangular chain mass matrix (Sec. IV.5); they do not rely on $\kappa_{a}\epsilon^{h_{a}/9}\ll 1$. The CKM matrix arises from the mismatch between the left-handed rotations $U_{dL}$ and $U_{uL}$ that diagonalize the two Yukawa textures.
@@ VI.7 Jarlskog invariant scaling
$J\sim\epsilon^{8/9}\cdot\epsilon^{17/9}\cdot\epsilon^{10/3}\cdot\sin\delta\sim\epsilon^{55/9}\,\sin\delta\simeq 3.5\times 10^{-5}\sin\delta,$ (51)
@@ VII.2 Genericity of $\mathcal{O}(1)$ CP violation
Phase counting. The standard phase-counting theorem for three generations gives (n1)(n2)/2=1(n-1)(n-2)/2=1 physical CP-violating phase [24]. The multi-messenger phases $\phi_{f}$, $\psi_{f}$ are not additional parameters: they are specific combinations of the UV coupling phases that survive after all field rephasing. The chain framework therefore has the same number of irreducible phase inputs as the SM—one. Its advantage is that the lattice exponent algebra determines the scaling $J\sim\epsilon^{55/9}\sin\delta$ structurally, while the dense ninths spacing ensures that the remaining free parameter $\sin\delta$ is generically $\mathcal{O}(1)$.
@@ XII.1 The charged-lepton GMR and $\epsilon^{3/2}$
and numerically $\epsilon^{3/2}\simeq 0.0806$—an exact match. The exponent $3/2=Q(L_{1})+Q(L_{2})=1+\tfrac{1}{2}$ is the sum of the first- and second-generation lepton-doublet charges, so $\beta$ is determined by the $\mathbb{Z}_{18}$ discrete gauge charges.
@@ XII.4 Dual mass representations of CKM elements
These identities are not numerological coincidences: they are algebraic consequences of the fact that all six quark masses are determined by a single parameter $B$ and the discrete $\mathbb{Z}_{9}$ charges. In the $B$-lattice, the masses scale as $(u,c,t)=t\,(\epsilon^{64/9},\,\epsilon^{10/3},\,1)$ and $(d,s,b)=b\,(\epsilon^{37/9},\,\epsilon^{7/3},\,1)$, so any ratio of mass monomials reduces to a rational power of $\epsilon$. A complete catalog of the resulting dual CKM representations and geometric mean relations will be presented in a forthcoming companion paper.
@@ XIII Discrete Anomaly Cancellation
XIII.1 $\mathbb{Z}_{18}$ charge assignments
The complete $\mathbb{Z}_{18}$ charge table follows from the $B$-lattice FN charges determined in Secs. V and XI [5, 21], expressed as integer charges $q_{18}\equiv 18\,Q$:
@@ XIII.1 $\mathbb{Z}_{18}$ charge assignments
Here $N_{i}$ are right-handed (sterile) neutrinos and $H$ is the SM Higgs doublet. The third generation of all SM fields carries zero charge, so the anomaly sums receive contributions only from the first two generations. The $\mathbb{Z}_{9}$ subgroup (denominator-$9$ lattice) controls quark-sector exponents, while the full $\mathbb{Z}_{18}$ is required to accommodate the half-integer lepton-doublet charge $Q(L_{2})=\tfrac{1}{2}$.
@@ XIII.2 Neutrino masses from the charge table
The $\mathbb{Z}_{18}$ charges determine the neutrino mass texture through two complementary mechanisms.
@@
The dimension-five operator $L_{i}L_{j}HH/\Lambda$ carries $\mathbb{Z}_{18}$ charge $Q(L_{i})+Q(L_{j})$, so the required number of flavon insertions gives the exponent matrix $p^{\nu}_{ij}=Q(L_{i})+Q(L_{j})$ of Eq. (79).
@@
The vanishing $\mathbb{Z}_{18}$ charge of $N_{i}$ has two further consequences: the right-handed neutrinos are SM singlets with zero discrete charge, so they do not contribute to any of the anomaly sums (106)–(108); and their Majorana mass $M_{N}\sim\Lambda$ is unsuppressed by the flavon, placing the seesaw scale at $\Lambda\sim 5\times 10^{11}$ GeV, consistent with the axion decay constant $f_{a}$ discussed in Sec. XIV.
@@ XIII.3 Green–Schwarz cancellation and the axion
The SM fermion charges above do not by themselves satisfy Eqs. (106)–(108); the residual anomalies $A_{a}\not\equiv 0\pmod{18}$ are nonzero. This is a generic feature of FN models and is resolved by the Green–Schwarz (GS) mechanism: the $\mathbb{Z}_{18}$ discrete gauge symmetry descends from a continuous $U(1)_{\rm FN}$ that is spontaneously broken by the flavon VEV $\langle\Phi\rangle$. The $U(1)_{\rm FN}$ mixed anomalies are cancelled by a GS axion, and this cancellation persists in the discrete $Z_{N}$ remnant [4, 38].
The connection to the strong CP problem is immediate: the same PQ axion that solves the strong CP problem (Sec. XIV) can serve as the GS field that cancels the discrete gauge anomalies. This identification is natural in the “Flavor in Ninths” framework [21], where the $\mathbb{Z}_{18}$ discrete gauge symmetry simultaneously
@@
The VLQ chain fields ($D_{a}+\bar{D}_{a}$, $U_{a}+\bar{U}_{a}$) are vectorlike under all gauge groups including $\mathbb{Z}_{18}$ and therefore do not contribute to the anomaly sums.
@@ XIII.4 Spectator fermion alternative
If the GS mechanism is not invoked, anomaly cancellation can alternatively be achieved by introducing spectator fermions at a high scale $\Lambda\gg M_{\rm VLQ}$ that carry appropriate $\mathbb{Z}_{18}$ charges but are SM singlets. Such spectators are common in UV completions of FN models and decouple from low-energy physics. In either case, the anomaly-free $\mathbb{Z}_{18}$ is a legitimate discrete gauge symmetry, and its protection of the axion quality and the lattice exponent structure is on firm theoretical footing.
@@ XIV Connection to the Axion Quality Problem
The “Flavor in Ninths” paper [21] demonstrated that the same discrete $Z_{N}$ symmetry enforcing the ninths lattice can simultaneously address the Peccei–Quinn (PQ) axion quality problem [39]. The axion quality problem—the requirement that Planck-suppressed operators violating the global $U(1)_{\rm PQ}$ symmetry be forbidden to sufficiently high dimension—was identified by Kamionkowski and March-Russell [40], Holman et al. [41], and Barr and Seckel [42], and has been studied quantitatively in the context of the string landscape by Baer, Barger, and Sengupta [43] and in gauge-protected axion models by Bhattiprolu and Martin [44]. In the UF construction, the FN flavon $\Phi$ is identified with the PQ scalar, and the $\mathbb{Z}_{18}$ discrete gauge symmetry forbids all PQ-violating Planck-suppressed operators below dimension $9$.
@@
This connection carries over to the chain framework: the $\mathbb{Z}_{9}$ (or $\mathbb{Z}_{18}$) symmetry that enforces nearest-neighbor locality along the VLQ chain is the same symmetry that protects the axion potential. The chain construction therefore inherits the axion-quality protection automatically, without requiring any additional symmetry.
@@
The axion in this mass window ($f_{a}\sim 10^{11}$ GeV) is a viable dark matter candidate. The continued absence of weakly interacting massive particle (WIMP) signals in direct detection experiments [47, 48, 49] suggests that WIMPs—if they were ever thermally produced—may have decayed away via small symmetry-violating couplings, leaving axions as the dominant component of dark matter today [50, 51]. In the UF framework this outcome is natural: the same $\mathbb{Z}_{18}$ discrete gauge symmetry that protects the axion quality also forbids the lowest-dimension operators that would destabilize the lightest supersymmetric particle in SUSY extensions, providing a unified explanation for both the flavor hierarchy and the dark matter composition.
@@ XVI Discussion
    The discrete anomaly of $\mathbb{Z}_{18}$ is cancelled by the Green–Schwarz mechanism, identifying the GS axion with the PQ axion that solves the strong CP problem. The specific group order $N=18$ is not a free choice: it is the minimal $Z_{N}$ that simultaneously accommodates the half-integer lepton doublet charge $Q(L_{2})=\tfrac{1}{2}$, the ninths-quantized quark exponents, and the axion quality requirement that PQ-violating operators be forbidden below dimension $N/2=9$.
@@
Axion quality $\mathbb{Z}_{18}$ protected Haloscopes
@@ XVII Conclusions
Unified Flavor provides the dynamical ultraviolet completion of the $B$-lattice hierarchy program. A single flavon $\Phi$ with $\epsilon=1/B=14/75$ enforces discrete ninths-quantized Yukawa exponents, while TeV-scale nearest-neighbor VLQ chains generate the effective Yukawa couplings as algebraic path sums. Each effective Yukawa amplitude factorizes into entry, chain-propagation, and exit factors controlled by the discrete $\mathbb{Z}_{9}$ charges of the SM fields, with the mass hierarchy encoded in the exponents $\epsilon^{Q(Q_{i})+Q(f^{c}_{j})}$. A multi-messenger structure—coherent contributions from several chain configurations—generates $\mathcal{O}(1)$ complex coefficients whose phases are the physical origin of CKM CP violation.
@@
Unified Flavor completes the $B$-lattice series by demonstrating that the phenomenologically successful lattice exponent algebra emerges from a definite symmetry structure and field content, rendering the fermion mass hierarchy a structural consequence of the UV theory rather than a parametric accident. For convenience, Appendix E collects the two calculation rules (Suppression Rule and Hops Rule) together with comprehensive reference tables for all Yukawa suppressions and their chain-diagram decompositions.
@@ D.3 Physical interpretation
(iii) One irreducible phase. The standard phase-counting theorem gives (n1)(n2)/2=1(n-1)(n-2)/2=1 physical CP phase for three generations. The multi-messenger phases $\phi_{f}$, $\psi_{f}$ are specific combinations of the UV Yukawa coupling phases that survive after all field rephasing; they do not represent additional parametric freedom. The framework’s structural claim is that the Jarlskog scaling $J\sim\epsilon^{55/9}\sin\delta$ is fixed by the lattice exponents, while the remaining free parameter $\sin\delta$ is generically $\mathcal{O}(1)$—it is not tuned to be either large or small.